Shift report 20010803 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Aug 03 2001 - 16:29:58 EDT

  • Next message: Ion-Sorin Zgura: "Re: Urgent!! Final draft of multiplicity paper needs comments NOW!"

    ------------------------------------------------------------
    Date: Fri Aug  3 16:29:57 2001, Shift: 08:00-16:00
    Supervisor: CR, with: 
    Log book#7 pages 16 - 16
    
    Shift Summary:
    ------------------------------------------------------------
    1. All the shift there were devlopements for RF, and   
       some technical difficulties. 
    2. #3986 - Pedestal run for all detectors.
       #3987 - Pedestal run TPCs
    3. Chellis and Hiro fixed T1. See further instructions in 
       logbook 7 page 16.
    DB:
     - Probably No Beam for 24~48 hrs.   
    	* Fire in Main Motor generator
    	* AGS is switching to alternate generator will take
    	  12~48 hrs to complete.	
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Aug 03 2001 - 16:30:58 EDT